OpenStudy (anonymous):

in how many three digits number can be formed if ____a.there are no restrictions? ____b.all three digits is greater than 700? ____c.the number is less than 200? ____d.the number is greater than 700? ____e.the number is greater than 600 but greater than 400? ____f.the number is greater than 100,divisible by 5 and no two digits are the same? ____g.the number is less than 800 but greater than 400,is even and no two digits are repeated?

OpenStudy (anonymous):

no restrictions? 1000

OpenStudy (anonymous):

second one doesn't make any sense

OpenStudy (akshay_budhkar):

second 1 300?

OpenStudy (anonymous):

number is less than 200 first digit is oh wait. three digit number. ok so least it can be is 100

OpenStudy (anonymous):

most it can be is 999

OpenStudy (amistre64):

that is alot of questions to answer for free/one medal :)

OpenStudy (anonymous):

there are 900 of them

OpenStudy (anonymous):

yes no restriction sir

OpenStudy (akshay_budhkar):

was that for me amistre??lol

OpenStudy (anonymous):

now if it is less than 200 first digit is a 1 so you have 10 times 10 = 100 choices for last two

OpenStudy (anonymous):

a. 810 b. 8 c. 100 d. 200 e. wrong question f. 128 g. 120

OpenStudy (anonymous):

well "no restrictions" is fine, but i don't call 002 a three digit number

OpenStudy (akshay_budhkar):

ramkrishna???

OpenStudy (anonymous):

if you think 002 counts as a three digit number, then there are 1000 of them, 000 through 999

OpenStudy (anonymous):

002 is one digit number not 3 digit number

OpenStudy (anonymous):

where is the 002?

OpenStudy (akshay_budhkar):

i believe ur answes r wrong ramkrishna?

OpenStudy (anonymous):

"satellite73" is talking about 002

OpenStudy (anonymous):

ramkrishma e. question >> the number is less than 600 but greater than 400?

OpenStudy (anonymous):

Akshay Budhkar all the answers are correct . tell me if any particular answer is wrong and why?

OpenStudy (anonymous):

answer of part e. 100

OpenStudy (anonymous):

if the first digit has to be at least a 1, then doesnt that make 900 three digit numbers? (for no restrictions) 9 possible numbers for first digit 10 for second 10 for third Multiply and get 900? Or do 999-100+1 = 900?

OpenStudy (anonymous):

or am i reading that wrong.

OpenStudy (anonymous):

lets do f) greater than 100 means the first digit has 8 choices divisible by 5 means last digit has two choices, 0 or 5 no two digits are the same means ... good luck!

OpenStudy (akshay_budhkar):

b is wrong ramkrishna

OpenStudy (anonymous):

love to see a clear answer to f) but i am going for lunch. yikes

OpenStudy (anonymous):

ok ok guys let see this example of my prof > how many three digitscan be formed from the numbers 1,2,3,4,5 ____a.repetition is allowed? ____b.repetition is not allowed? ok<<<<this is the example

OpenStudy (anonymous):

answer given by me for part a was wrong correct answer a. 900

OpenStudy (anonymous):

try to answer the example of my prof

OpenStudy (anonymous):

explanation for b) since all the three digits should be greater than 7 so the possibilty of all the three places is 2. so. 2*2*2=8

OpenStudy (anonymous):

correct sir

OpenStudy (anonymous):

if repetition is allowed, we have 5 possible choices for first digit 5 for second 5 for 3rd 5 for 4th and 5 for 5th. 5*5*5*5*5 = 3125

OpenStudy (anonymous):

how many three digitscan be formed from the numbers 1,2,3,4,5 ____a.repetition is allowed? ____b.repetition is not allowed? ans. a) 125 b) 60

OpenStudy (anonymous):

i thought 5= (5)(4)(3)(2)(1)= 120

OpenStudy (anonymous):

if repeatition is not allowed

OpenStudy (anonymous):

if repetition isnt allowed we have: 5 choices for 1st number 4 for second 3 for third, 2 for 4th 1 for 5th. thats 5! which is 120

OpenStudy (anonymous):

im sorry guys got confused letter b question is here ____b.all three digits must be the same?

OpenStudy (anonymous):

wow im just reading what i want to =/

OpenStudy (anonymous):

ramkrishna is correct

OpenStudy (anonymous):

b) 9

OpenStudy (anonymous):

somehow i was thinking we needed a 5 digit number

OpenStudy (anonymous):

if all three digits must be same then there are 9 possibility for the first place then the possibility for the remaning two places is 1. so 9*1*1=9

OpenStudy (anonymous):

if you need explanation for any part then tell me.

OpenStudy (anonymous):

joemath correct in repetion is not allowed

OpenStudy (zarkon):

isn't (f) ... 135

OpenStudy (anonymous):

How? according to me since the number should be greater than 100 so first digit can be arranged in 8 possible ways. and it should be divisible by 5 so the unit place should be either 0 or 5. and now we left with only 8 digits since the repetition is not allowed so 8*8*2=128

OpenStudy (zarkon):

That doesn't work

OpenStudy (zarkon):

if we didn't care about the number being greater than 100 we could do this 2 choices for the ones digit, 9 choices for the 10's digit and 8 choices for the 100's digit that gives 2*9*8=144 possibilities included in that are things we don't want like 100 so we are down to 143 choices finally any number of the form 0X5 (0 in the 100's position and 5 in the 1's position) is not wanted ...there are 8 of these 143-8 135